Grade 12th passAlgebra

Write a Pythagorean triplet whose one member is. ii. 14

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

2m=14⇒m=14/2=7

m2–1= 72–1 = 49–1 = 48

m2+1=72+1=49+1=50
∴ (14, 48, 50) is not a Pythagorean triplet.
Thanks